Find the perpendicular distance of the point (2, -1) from the line y= 3x+1
Answer by Alan3354(69443)

Find the perpendicular distance of the point (2, -1) from the line y= 3x+1
-------------
Find the perpendicular distance of the point (2,-1) from the line
3x - y + 1 = 0
a = 3
b = -1
c = 1
----
d = (ax + by + c)/sqrt(a^2 + b^2)
d = (3*2 + 1 + 1)/sqrt(9+1) = 8/sqrt(10)
d = 0.8*sqrt(10)